OpenStudy (anonymous):

Given f(x) = 5x - 6 and g(x) = x2 - 5x + 6, find the function (f - g)(x). Use the ^ key for the exponent. Will someone please help me with this and explain how to get the answer?!..

jimthompson5910 (jim_thompson5910):

(f - g)(x) = f(x) - g(x) (f - g)(x) = (5x-6) - (x^2-5x+6) From here, distribute and combine like terms

OpenStudy (anonymous):

so it would be x^2+6?

jimthompson5910 (jim_thompson5910):

(f - g)(x) = (5x-6) - (x^2-5x+6) (f - g)(x) = 5x-6 - x^2-5x+6 (f - g)(x) = 5x-6 - x^2+5x-6 What's next?

OpenStudy (anonymous):

cyou would cancel out the 5x's right?

jimthompson5910 (jim_thompson5910):

why would they cancel?

OpenStudy (anonymous):

cause 5x-5x? I'm not sure this is really confusing me..

jimthompson5910 (jim_thompson5910):

no the negative distributes and changes it to +5x

OpenStudy (anonymous):

oh so then would it be 10x^2-12

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

(f - g)(x) = 5x-6 - x^2+5x-6 (f - g)(x) =-x^2 + (5x+5x)+ (-6-6) (f - g)(x) = -x^2 + 10x - 12

OpenStudy (anonymous):

oh.. I see how you broke it down, I kept canceling them out and i was left with just the x^2 thank you!

jimthompson5910 (jim_thompson5910):

np